UT 20131218 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, MEdwards
Telescope Operator: DGHuerta
LBTO Support:

Plan for the night

- MOS, MOS, MOS-acquisition
- Functional checkout AdSec right
- Verification new AGw-transform
- Determination of center of rotation

Summary

- MOS-acqusition finally works!

- AdSec checkout successfull

- NewAGw -transform in place

- Image quality, image scale and distortion

- Determination of center of rotation

- Accuracy of offsets

- Half a night with 0.3" on guider smile and 2-3mag extinction due to Cirrus frown, sad smile

Preparations

- Ingo provided a new version of his tool, implemented by Alexander

Details

All times are UT, files are luci2.20131217.iiii

20:04 #1-10, no idea who has taken the data, probably darks

23:05 #11, ??

23:29, #12-19, 600sec darks, LIR, show persistence effects from earlier images

00:20 Sunset, quite many cirrus around, low wind, T=7deg, hum=15%, no flats

00:40 David checks pointing and collimates

MOS acquisition

01:03 Preset to M225UD3H1, N3.75, more clouds, thicker, takes a little while to collimate
#20 sky, #21 source, #22 MOS, aligned, offset dx=1.65, dy=-4", ddelta=2deg numbers differently from what we got last night), verification throughslit #23, 6 stars are in slit, some slight offset to dy necessary, distortion apparent, but no extra offsaet necessary to put stars in the box, verification wo MOS #24

Red crosses which indicate positions and real positions of stars are systematically
offset => need to recenter all 4 stars manually

MOS-alignment with N3.75 works

01:44 Redo with N1.8, #25 sky, #26 src, #27 MOS (N3.75), #28 MOS (N1.8), aligned, dx=2.4", dy"-3.7", ddelta=1.7deg, verification#29, wo MOS #30, do not see anything, redid verification with MOS (this time 20*2.75sec integrated and not simply 5sec) #31, OK

MOS-alignment with N1.8 works

Seeing about 0.5"

02:35 Redo Preset to M225UD3H1, modified Jarons script that it works
Michelle does the alignment and than we will try to take some spectra
useful to compare data quality to the one of LUCI1
sky #32-37, src #38-42, MOS #43, set to normal, revised script and restart,
sky #44, scr # 45, MOS #46, script went to 4th obsitem (not clear if this was human
or error instrinsic to script, #46 crap
Redo, #47 sky, #48 src, script hangs after 2nd obsitem, executed instr. setup., but didn't
take an image, restart SW, didn't want to collimate,
#49 MOS without collimation, moon to close (60deg away)??
David is pointing to a field far away to check GCS/TCS, OK,
we can not take spectra here frown, sad smile

03:51 Preset to M225UD3H2, new MOS-alignment, #51 sky, #52 src, #53 MOS
Michelle aligns again, verification #54, works (Michelle doesn't believe it)

4:45 Preset to M225UD3H2, new MOS-alignment, need to correct pointing,
need to collimate, guide star in Jarons script is galaxy...need to change that appropriately

05:00 Handover to JHill, DMiller for AdSec checkout

07:xx Starting AGw-transform verification

Image quality is exiting, down to 0.3" on guider !!!

07:35 Done with AdSec checkout and AGw-transform verification (Michelle starts to work on the latter)

Image quality N3.75 camera

Seeing is so good that we take data of Palomar 2 in zJHKKs
Need to find out if good astrometry of this globular cluster exists (should be)
if so, data can be used for image scale and distortion determination

07:55 Palomar2_abs_all, Ks #93-102, IQ ~3pix, J #103-112, IQ ~ 3pix,
Cirrus passing through, H #113, IQ ~ 3pix, script stopped after 1st H-band expoure,
gnagnagna, GEIRS socket is busy, Alexander do not know why
re-ini ROE, shutdown GEIRS, rstart GEIRS, etc..., #114 crap, restartet with
H #115-124 IQ~ 3.2pix, K #125-134 IQ ~ 3.5pix, z #135-144 IQ ~ 3.6 -> 5 pix
Extinction due to cirrus at least 3mag, surprising how well guider works even when moon and
cirrus illuminate the spots but seeing is good

Center of rotation

09:47 BS9126, PA=0, #145-168, yak, images not integrated, again for homogeneity,
#169/170, PA=180, #171/172, Pa=90, script hangs after preset, again GEIRS socket,
fixed by Alexander, #173 crap, #174/175, Pa=270, #176/177

MOS-acquisition

10:46 SDSS1110, Jarons script modified by Michelle, #178 sky, #179 src, #180 MOS,
alignment, dx=0.008", dy=-3.2", ddelta=1.9deg, verification #181, OK,
wo MOS #182, star in slit #33 very weak, hard to see with 3mag extinction,
taking spectra 2 * 300sec just to see if there is something Ks150 (stabilizes),
#183/#184, barely something, 3mag extinction is too much...

11:25 Telluric for SDSS1110, script hangs again, Alexander fixed it,
resent preset, no GS, correct pointing, OK, #185 crap, #186 sky,
#187 src, #188 MOS, no slits (HeI /z), #189 with z #190, aligned,
#191 OK, spectra #192-197

Tests - offsets

12:20 SDSS1245, J, #198-203, stopped after 6th exp., too many script hickups tonight,
Alexander decides to reboot Luci-machine, took a while

13:17 Restart SDSS1245, Ks, #204-225 IQ ~ 3pix,

13:50 Stopped, sky too bright, no flats due to Cirrus

13:55 Closing

14:42 Dark-script started, 10*1800sec, MER, #226-235,
10*(20*5sec int), LIR, #236-245,
10*(5*12sec int), LIR, #246-255,
10*(6*10sec int), LIR, #256-265,

-- JochenHeidt - 18 Dec 2013
Topic revision: r11 - 04 Jun 2014, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback